OpenStudy (anonymous):

solve the matrix equation PLEASE HELP 5x+6y=14 4x-y=17

OpenStudy (nubeer):

familiar with cramers rule?

OpenStudy (anonymous):

no kinda can u help

OpenStudy (nubeer):

\[\left[\begin{matrix}5 & 6\\ 4 & -1\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right) = \left(\begin{matrix}14 \\ 17\end{matrix}\right)\] AX=B

OpenStudy (anonymous):

can u explain?

OpenStudy (nubeer):

can you find the determinent of A ?

OpenStudy (anonymous):

no :(

OpenStudy (nubeer):

you dont know how to find determinent? and u have to solve this through matrix? are you sure about it?

terenzreignz (terenzreignz):

Well, to make it brief, to get the determinant of a 2x2 matrix \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] its determinant is given by ad - bc

terenzreignz (terenzreignz):

Okay? So for instance, if we have the matrix \[\left[\begin{matrix}2 & 6 \\ -1 & 5\end{matrix}\right]\] Its determinant is (2)(5) - (-1)(6) = 16

OpenStudy (anonymous):

so what do u get?

terenzreignz (terenzreignz):

hold on :D Let's just put that equation up again :D 5x + 6y = 14 4x - y = 17 Now, construct the matrix where its components are the respective coefficients of x and y. (I realise that's vague, so here's how to do it) If your equations are of the form ax + by = h mx + ny = k Then your matrix A is of the form \[A = \left[\begin{matrix}a & b \\ m & n\end{matrix}\right]\]

terenzreignz (terenzreignz):

All right, I'll do that then... But please find its determinant, we need it for future use. Your matrix A is given by \[A = \left[\begin{matrix}5 & 6 \\ 4 & -1\end{matrix}\right]\] any questions? If none, then please proceed to finding its determinant :D

OpenStudy (anonymous):

okay more help?

terenzreignz (terenzreignz):

Formula for determinant was given above... please refer to it :D I also gave an example. Here's a refresher For a matrix M \[M=\left[\begin{matrix}w & x \\ y & z\end{matrix}\right]\] Its determinant is given by this formula : wz - yx

OpenStudy (anonymous):

okay so would u get 4 and -1

terenzreignz (terenzreignz):

Those are the coefficients of x and y respectively in the second equation, right? I put them in their proper places in the matrix A, as shown above. Could you find the determinant of A?

OpenStudy (anonymous):

no help

OpenStudy (anonymous):

@bri4life14 you're excellent of saying "NO" !!!

terenzreignz (terenzreignz):

Again, here's the sample matrix \[\left[\begin{matrix}2 & 6 \\ -1 & 5\end{matrix}\right]\] You multiply the upper left component to the lower right 2 x 5 you get 10 From that, you subtract the product of the lower left component and the upper right 10 - (-1)(6) 10 + 6 16 Get it?

OpenStudy (anonymous):

thanks

terenzreignz (terenzreignz):

Hey, @bri4life14 That was just an EXAMPLE I'd still prefer you evaluate the determinant of \[A = \left[\begin{matrix}5 & 6 \\ 4 & -1\end{matrix}\right]\]on your own, I know you can do it :)

OpenStudy (anonymous):

oh yea i know thanks @terenzreignz you have been very helpful :) but @Chlorophyll your rude and you can keep your comments to your self... if your not going to help get off my problem!